OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 8) and a directrix at y = -8.

hero (hero):

You're given two points: the focus (0,8) and directix (x,-8). Insert them in to this formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) Then simplify

OpenStudy (anonymous):

is it y2=8x @Hero ?

hero (hero):

http://sketchtoy.com/62817481

OpenStudy (anonymous):

thank you so much! that website was pretty cool too! @Hero

hero (hero):

Did you understand every step?

OpenStudy (anonymous):

yes i see what i did wrong that was helpful

hero (hero):

What rule did I use to get from \((y + 8)^2 - (y - 8)^2\) to \((y + 8 + y - 8)(y + 8 - (y - 8))\) ?

OpenStudy (anonymous):

didnt you just expand?

hero (hero):

Nope. I used Difference of Squares: \(a^2 - b^2 = (a + b)(a - b)\)